2. ANOVA problem

Chobani Greek Yogurt, which advertises its product as “100% natural, no artificial ingredients,” recently recalled its product after its cups were found to have mold. After the quality issue was diagnosed and rectified, Chobani’s marketing department took over, launching a media blitz to address consumers’ concerns with full-page ads in People magazine and the New York Times, distribution of free samples, and a social media campaign. Social Marketing Director Alison Chanes has devised five possible tweets, and has test marketed each message with 100 unique, carefully-selected members of the target market: 100 see message one, 100 see message two, etc. – so she recruited 500 participants overall. Each participant is asked to rate the given tweeted message on a 1-to-5 scale, with 5=best. Messages, with their mean and standard deviation ratings are given in the table below.

Message

n

mean

Standard dev

#ChotallyAwesome!

100

4.36

1.14

#MoldIsNatural!

100

2.89

0.87

#NewFlavorsAllTheTime

100

1.55

0.99

#FreePenicillin

100

3.17

1.36

#CulturallyDeprived

100

2.24

0.53

Alison has begun the analysis of the message test, but her virus software expired and her computer shut down before she could retrieve the final answer. But she did get sums of squares: SSE= 513.9, and SSA=444.6. Determine whether there’s a better message @ 95% confidence, and show all work.
Answer #1

k = 5

N = 500

df Between = k-1

df Total = N-1

MS = SS/df

F= MS /MS Error

p-value = 0.0000

since p-value < alpha, we reject the null hypothesis

we conclude that there is significant evidence of difference

Please give me a thumbs-up if this helps you out. Thank you! :)